OpenStudy (chaylaceyx3):

if z=a+bi and -z=a-bi, solve z+6(-z)=7 for z

OpenStudy (michele_laino):

hint: if z= a+ bi, then: -z = -a - bi

OpenStudy (michele_laino):

furthermore: z-6z=-5z

OpenStudy (chaylaceyx3):

sorry its actually the conjugate of z, not -z so its a-bi

OpenStudy (michele_laino):

ok! then your equation can be rewritten as follows: \[\Large z + 6\bar z = a + bi + 6a - 6bi = 7a - 5bi = 7\]

OpenStudy (michele_laino):

which is equivalent to these two real equations: \[\Large \left\{ \begin{gathered} 7a = 7 \hfill \\ - 5b = 0 \hfill \\ \end{gathered} \right.\] please solve that system for a and b

OpenStudy (chaylaceyx3):

a=1 and b=0 ?

OpenStudy (michele_laino):

that's right! so, what is z=...?

OpenStudy (chaylaceyx3):

do i just plug these numbers into 7a-5bi?

OpenStudy (michele_laino):

no, you have to plug, those values, for a and b, into z=a+bi

OpenStudy (michele_laino):

what do you get?

OpenStudy (chaylaceyx3):

so its 1?

OpenStudy (michele_laino):

correct! z=1 is the solution of your complex equation

OpenStudy (chaylaceyx3):

OH ok! Thanks so much!

OpenStudy (michele_laino):

:)

OpenStudy (chaylaceyx3):

question, where does the i go? my teacher asked and i wasn't sure